Solve the given bernoulli equation


dy/dx=y(xy^3 -1)


1
Expert's answer
2020-08-25T13:01:12-0400

dy/dx = y(xy3-1)=> dy/dx = xy4-y=> dy/dx + y = xy4 Dividing the above equation by y3 to get -


y-4 (dy/dx) + y-3 = x ---- A


Put y-3 = t ,then -3y-4 (dy/dx) =(dt/dx)
=> y-4(dy/dx) = (-1/3) (dt/dx)
Substituting above in equation A to get -(-1/3) (dt/dx) + t = x Multiplying both sides by -3=> dt/dx +(-3)t = -3xThe above resides as a linear first order equation of the form-t'+P(x)t=Q(x), where P(x)=-3 and Q(x)=-3xSo, the integrating factor for above bernoulli equation is - M(x)=exp( "\int"-3 dx) = e-3x.Multiplying this by both sides of the equation -
e-3x t' + (-3)t e-3x =-3x e-3xAs left hand side represents d/dx(t e-3x) :d/dx(t e-3x) = -3xe-3x Integrating both sides with respect to x :"\int" d/dx(t e-3x) dx = "\int" -3xe-3x dx=> t e-3x = (-3) [ x"\int" e-3x dx - "\int" {{(d/dx) x} "\int" e-3x dx} ]"\implies"t e-3x = (-3) [ (-1/3) x e-3x - "\int" 1. (-1/3) e-3x dx ]"\implies"t e-3x = x e-3x + (-1/3) e-3x "\implies"t = x -(1/3) Put t = y-3 to get - y-3 = x - (1/3)